OpenStudy (anonymous):

At the beginning of the year, the price of stock A was 35% greater than the price of stock B. Over the course of the year, the price of stock A doubles while the price of stock B decreased by 10%. At the end of the year, how many times greater was the price of stock A than the price of stock B?

OpenStudy (tkhunny):

Beginning: \(\dfrac{A}{B} = 1.35\) End: \(\dfrac{2A}{0.9B} = \dfrac{2}{0.9}\cdot\dfrac{A}{B}\) Now what?